Circular No. 9089

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V5584 SAGITTARII = NOVA SAGITTARII 2009 No. 4 Koichi Nishiyama, Kurume, Fukuoka-ken, Japan; and F. Kabashima,

Miyaki-cho, Saga-ken, Japan, report their discovery of an apparent nova (mag 9.3) on two 60-s unfiltered CCD frames (limiting mag 13.6) taken on Oct. 26.439 and 26.440 UT using a 105-mm f/4 camera lens; their five unfiltered CCD frames (limiting mag 17.5) taken around Oct. 26.455 using a 0.40-m reflector yield mag 9.3 and the following position for the new star: α = 18^{h}31^{m}32^{s}.79, δ = −16^{o}19′07′′.5 (equinox 2000.0). Nothing is visible at this position on their recent survey frames taken on Oct. 20.449 (limiting mag 13.9) and 21.451 (limiting mag 13.4). Following posting on the Central Bureau’s unconfirmed-objects webpage, other observers have reported CCD magnitudes and astrometry: Oct. 26.764, mag 9.3, position end figures 32^{s}.79, 07′′.8 (P. Corelli, Pagnacco, Italy, 0.45-m reflector); Oct. 27.09, ∼ 9.0, 32^{s}.81, 07′′.5 (E. Guido and G. Sostero, remotely with a 0.25-m reflector at the GRAS Observatory near Mayhill, NM, U.S.A.). Corelli adds that nothing is visible at the nova’s position on a Palomar plate (limiting mag 21.0). Additional details appear on CBETs 1994 and 1995.

K. Kinugasa, S. Honda, O. Hashimoto, H. Taguchi and H. Takahashi, Gunma Astronomical Observatory (GAO), write that a low-resolution spectrum (range 420-800 nm; resolution 1.0 nm), obtained with the GAO 1.5-m telescope (+ GLOWS) on Oct. 27.4 UT, confirms that this variable star is a nova, showing Hα emission (FWHM ∼ 600 km/s) with a P-Cyg profile, its absorption minimum being blue-shifted by 900 km/s with respect to the emission peak. Fe II (multiplet 42) lines also have P-Cyg profiles.

H. Maehara, Kwasan Observatory, Kyoto University, writes that a lowresolution optical spectrum (range 400–800 nm, resolution ∼ 200) of this nova was obtained on Oct. 27.42 UT with the 25-cm telescope of the Kwasan Observatory. The spectrum shows an H-alpha emission line that suggests that this object is a classical nova. CCD images of the nova yield the following magnitudes: Oct. 27.460, I_{c} = 8.66; 27.461, R_{c} = 9.29; 27.461, V = 9.90; 27.463, B = 11.06. Maehara also forwards the report of a spectrogram obtained by M. Fujii (Okayama, Japan; 40-cm telescope, range 400–950 nm) on Oct. 27.43, showing Hα, Hβ, and Fe II lines with P-Cyg features; the absorption component of the Hα line is blue-shifted from the emission peak by 1.5 nm. These features suggest that this object is a classical nova around maximum. Fujii adds that the Na D line has a equivalent width of 0.7 nm.

N. Samus and E. V. Kazarovets report that the GCVS team has assigned the designation V5584 Sgr to this nova.

2009 October 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green